# SMiR: Efficient Synthetic Data Pipeline to Improve Multi-image Reasoning

### v1

@@ 4 Multi-Image Benchmark
We create questions that challenge Vision Language Models (VLMs) to reason over multiple related images, analyze relationships, and derive meaning from series of images. The diverse range of tasks in this benchmark pushes the boundaries of multimodal understanding, requiring sophisticated analysis and interpretation across various visual contexts. Our benchmark compels models to provide detailed explanations for complex visual tasks, going beyond simple multiple-choice answers to demonstrate advanced reasoning capabilities. We implement an open-ended questionnaire with a judge model that evaluates responses in a pairwise structure, enabling relative comparisons between models and highlighting subtle differences in their ability to handle complex multi-image reasoning tasks.
@@ 4.1 Benchmark Overview
To address these challenges, we developed SMiR-Bench, a multi-turn benchmark consisting of 200 examples across seven diverse topics. These topics include:
@@
  • Matching: pairing photos based on visual cues
@@
  • Pattern: identifying visual task patterns
  • Ranking: ordering everyday objects based on preferences
  • Storytelling: narrating based on image sequences, as shown in Figure 4
  • Visual: finding connections between images
The benchmark was curated collaboratively by a human annotator and GPT-4, utilizing images from the internet and the Shot2Story dataset (Han et al., 2023).
@@ 4.2 Evaluation Methodology
SMiR-Bench extends the Auto-Hard-Auto v0.1 (Li et al., 2024c) framework to the multimodal domain. The evaluation process begins with all models completing the benchmark tasks. Subsequently, one model is selected as the baseline, and a GPT-4o judge conducts pairwise comparisons between the baseline model’s responses and those of every other model. For each comparison, the judge evaluates responses on helpfulness, relevance, and conciseness. The judge’s assessments result in comparative rankings between two models (represented as A and B), such as $A\gg B$ (A significantly better than B), $A>B$ (A better than B), $A=B$ (A equal to B), or their inverses. Based on these pairwise comparisons and confidence levels, the framework generates an overall ranking of all models. This comprehensive approach enables a multi-turn, automatic, and challenging evaluation process, providing a nuanced assessment of model performance in multimodal tasks. By leveraging pairwise comparisons and confidence levels, it offers a clear view of how models perform relative to each other, complete with statistical reliability measures.
@@
4.3 Results
To evaluate the effectiveness of our SMiR dataset, we finetune popular open-source VLMs and assess their performance on SMiR-Bench. Our experiments demonstrate significant improvements, with finetuned models achieving up to 8% better performance compared to their base versions.
4.4 Experiments
For our evaluation, we begin with two pre-trained models: Mantis-8B-siglip-llama3-pretrained (Jiang et al., 2024) and idefics-8b (Laurençon et al., 2024b). The first model combines a Llama-3-8B backbone with a SIGLIP encoder and is pretrained on CC3M (Gan et al., 2022). The second uses a Mistral-7B-v0.1 backbone with a perceiver (Jaegle et al., 2021) and MLP projection, pretrained on OBELICS. Both models were initially instruction-tuned on MANTIS-Instruct (721K examples) to create Mantis-8B-siglip-llama3 and Mantis-8B-Idefics2.
We then finetune these same pre-trained models on our SMiR dataset to create SMiR-8B-siglip-llama3-160 (trained on 160K examples) and SMiR-8B-Idefics2-100 (trained on 100K examples). To evaluate performance, we conduct pairwise comparisons using GPT-4 as a judge model against three specific baselines: GPT-4-Turbo to assess our models against a closed-source solution (Table 3), Mantis-8B-siglip-llama3 to evaluate SMiR-8B-siglip-llama3-160 (Table 4), and Mantis-8B-Idefics2 to compare with SMiR-8B-Idefics2-100 (Table 5). Remarkably, our SMiR-tuned models outperform their MANTIS-tuned counterparts despite being trained on a significantly smaller dataset. These results demonstrate the effectiveness of our SMiR dataset for finetuning, showing consistent performance improvements across all evaluation scenarios.
Table 3: SMiR-Bench Scores with GPT-4-Turbo Baseline
@@ 4.4 Experiments
SMiR-8B-siglip-llama3-160 5.0 +.8% (-2.2, 2.7) 183
@@
Table 4: SMiR-Bench Scores with Mantis-8B-siglip-llama3 Baseline
Model Name SMiR-Bench $\Delta$ 95% CI Average Tokens
Claude-3-Opus-20240229 97.4 (-1.3, 1.0) 321
Claude-3-5-Sonnet-20240620 97.1 (-1.3, 1.2) 362
GPT-4-Turbo 96.4 (-1.8, 1.0) 359
Gemini-1.5-Pro 96.3 (-1.3, 1.3) 361
GPT-4o 91.5 (-2.4, 1.8) 316
SMiR-8B-siglip-llama3-160 58.1 +8.1% (-4.6, 5.3) 156
Mantis-8B-siglip-llama3 50.0 (0.0, 0.0) 146
LLaVA-v1.6-Mistral-7B-HF 15.8 (-2.4, 3.1) 317
Mantis-8B-siglip-llama3-pretraind 8.7 (-2.1, 2.1) 207
Table 5: SMiR-Bench Scores with Mantis-8B-Idefics2 Baseline
Model Name SMiR-Bench $\Delta$ 95% CI Average Tokens
Claude-3-Opus-20240229 98.1 (-1.2, 0.9) 321
Claude-3-5-Sonnet-20240620 98.1 (-1.1, 1.1) 362
Gemini-1.5-Pro 97.5 (-1.1, 0.9) 361
GPT-4-Turbo 95.9 (-1.5, 1.6) 359
GPT-4o 94.5 (-1.6, 1.5) 316
SMiR-8B-Idefics2-100 58.0 +8.0% (-5.3, 3.9) 157
Mantis-8B-Idefics2 50.0 (0.0, 0.0) 171
Idefics2-8B 29.7 (-3.7, 4.4) 118
LLaVA-v1.6-Mistral-7B-HF 16.7 (-3.4, 3.0) 317
Figure 5: SMiR Dataset Size vs. Benchmark Score
5 Conclusion
This paper introduces a synthetic data pipeline designed to enhance multi-image reasoning capabilities on open-source VLMs. By leveraging multimodal embeddings and grouping algorithms, the pipeline generates high-quality synthetic multi-image reasoning instruction tuning data. The approach yields up to 8% improvement on SMiR-Bench for popular open-source models, demonstrating the significant potential of synthetic data in advancing open-source VLM models.
Limitations
Our methods have several limitations. Random sampling with iteration is time-intensive due to the need for recalculating distance embeddings for each new image sampled. Further investigation is needed to determine the scalability of our synthetic data. Future research should focus on developing more time-efficient algorithms and optimizing data mixtures.
Broader Impact
This paper introduces a method for generating high-quality, cost-effective data for VLMs, addressing the growing challenge of data scarcity. By advancing these open-source techniques, we contribute to narrowing the performance gap between open and closed-source models, promoting more accessible and powerful multimodal AI.
Appendix A Algorithm Details
A.1 Greedy Cluster Matching Algorithm
@@
A.2 Random Sampling with Iteration
@@ 4. Return $S$
Appendix B Data Samples
@@
B.1 Greedy Cluster Matching Algorithm
@@
B.2 Random Sampling with Iteration
@@ Appendix C Prompt
C.1 LLaVA Style Prompt

### v2

@@ 3 Multi-Image Benchmark
We introduce a benchmark that challenges VLMs to reason over multiple related images, analyze relationships, and derive meaning from image sequences. Unlike standard multiple-choice tasks, our benchmark requires detailed explanations, assessing models’ ability to perform complex multimodal reasoning. To enable robust comparisons, we employ an open-ended evaluation framework where a judge model ranks responses in a pairwise structure, highlighting subtle differences in reasoning capabilities.
@@ 3.1 Benchmark Overview
We introduce SMiR-Bench, a multi-turn benchmark consisting of 200 examples spanning seven diverse topics. These topics include:
@@
  • Matching: pairing photos based on visual similarities
@@
  • Pattern: recognizing visual patterns in structured tasks
  • Ranking: ordering objects based on contextual preferences
  • Storytelling: narrating events from image sequences, as shown in Fig. 3
  • Visual: establishing meaningful connections between images
For each category, we carefully selected 4-5 images from internet sources. Storytelling examples were specifically drawn from Shot2Story (Han et al., 2023) videos to leverage their sequential nature. GPT-4 was prompted to generate multi-turn question-answer conversations for these images, which were subsequently reviewed and refined by human annotators to ensure coherence and meaningful progression across dialogue turns.
@@ 3.2 Evaluation Methodology
SMiR-Bench extends the Auto-Hard-Auto v0.1 (Li et al., 2024b) framework to multimodal evaluation. The process begins with all models completing the benchmark tasks. One model is selected as the baseline, and a GPT-4o judge conducts pairwise comparisons between the baseline model’s responses and those of other models. The judge evaluates responses based on helpfulness, relevance, and conciseness, assigning comparative rankings such as $A\gg B$ (A significantly better than B), $A>B$ (A better than B), or $A=B$ (A equal to B). The framework aggregates these pairwise comparisons and confidence levels to produce an overall ranking of all models. An example of this evaluation process is illustrated in Fig. 3.
4 Results
To evaluate the effectiveness of our SMiR dataset, we finetune popular open-source VLMs and assess their performance on SMiR-Bench. Our experiments show that models fine-tuned on SMiR achieve up to an 8% improvement over their base versions, demonstrating the dataset’s value for enhancing multi-image reasoning capabilities.
We begin with two pre-trained models: Mantis-8B-siglip-llama3-pretrained (Jiang et al., 2024) and Idefics-8B (Laurençon et al., 2024b). The first model integrates a Llama-3-8B backbone with a SIGLIP encoder and is pre-trained on CC3M (Gan et al., 2022). The second employs a Mistral-7B-v0.1 backbone with a perceiver-based projection (Jaegle et al., 2021) and is pre-trained on OBELICS.
Our evaluation consists of multiple stages. First, both pre-trained models undergo instruction tuning on MANTIS-Instruct (721K examples), resulting in Mantis-8B-siglip-llama3 and Mantis-8B-Idefics2. We then fine-tune these same pre-trained models on our SMiR dataset, producing SMiR-8B-siglip-llama3 and SMiR-8B-Idefics2.
To assess model performance, we conduct pairwise comparisons using GPT-4o as a judge model against three baselines:
  • GPT-4-Turbo, to compare our models against a leading closed-source solution (Table 5).
  • Mantis-8B-siglip-llama3, to evaluate the improvements from SMiR fine-tuning (Table 3).
  • Mantis-8B-Idefics2, to assess the performance gains of the Idefics-based model trained on SMiR (Table 4).
Despite being trained on a significantly smaller dataset, models fine-tuned on SMiR outperform their MANTIS-tuned counterparts across all evaluation scenarios, demonstrating the value of our dataset for improving multi-image reasoning. Moreover, the significant performance gap between open-source and closed-source models on SMiR-Bench highlights the benchmark’s difficulty. Unlike prior benchmarks, SMiR-Bench focuses on highly correlated images, multi-turn question answering, and open-ended reasoning, making it a more rigorous test of multimodal understanding.
Table 3: SMiR-Bench scores with Mantis-8B-siglip-llama3 baseline.
Model Name SMiR-Bench $\Delta$ 95% CI Average Tokens
Claude-3-Opus-20240229 97.4 (-1.3, 1.0) 321
Claude-3-5-Sonnet-20240620 97.1 (-1.3, 1.2) 362
GPT-4-Turbo 96.4 (-1.8, 1.0) 359
Gemini-1.5-Pro 96.3 (-1.3, 1.3) 361
GPT-4o 91.5 (-2.4, 1.8) 316
SMiR-8B-siglip-llama3 58.1 +8.1% (-4.6, 5.3) 156
Mantis-8B-siglip-llama3 50.0 (0.0, 0.0) 146
LLaVA-v1.6-Mistral-7B-HF 15.8 (-2.4, 3.1) 317
Mantis-8B-siglip-llama3-pretraind 8.7 (-2.1, 2.1) 207
Table 4: SMiR-Bench scores with Mantis-8B-Idefics2 baseline.
Model Name SMiR-Bench $\Delta$ 95% CI Average Tokens
Claude-3-Opus-20240229 98.1 (-1.2, 0.9) 321
Claude-3-5-Sonnet-20240620 98.1 (-1.1, 1.1) 362
Gemini-1.5-Pro 97.5 (-1.1, 0.9) 361
GPT-4-Turbo 95.9 (-1.5, 1.6) 359
GPT-4o 94.5 (-1.6, 1.5) 316
SMiR-8B-Idefics2 58.0 +8.0% (-5.3, 3.9) 157
Mantis-8B-Idefics2 50.0 (0.0, 0.0) 171
Idefics2-8B 29.7 (-3.7, 4.4) 118
LLaVA-v1.6-Mistral-7B-HF 16.7 (-3.4, 3.0) 317
5 Conclusion
We introduces SMiR, a synthetic data pipeline for enhancing multi-image reasoning in open-source VLMs. By leveraging multimodal embeddings and grouping algorithms, our approach generates high-quality instruction tuning data, leading to up to an 8% improvement on SMiR-Bench.
However, certain limitations must be acknowledged. Random sampling with iteration is computationally intensive due to the need to recalculate distance embeddings for each newly sampled image. The scalability of our synthetic data also requires further investigation. Additionally, while SMiR-Bench provides a rigorous evaluation, our approach should be tested across other existing benchmarks to assess its generalizability.
Appendix A Related Works
Vision Language Models
We focus on instruction tuning Vision-Language Models (VLMs) that utilize a pretrained Large Language Model (LLM) backbone because this approach is cost-effective and more accessible for the open-source community. Since the backbone responsible for language understanding is already trained, the overall training process becomes simpler and requires fewer resources. Our primary task involves aligning the vision encoder—typically architectures like Vision Transformer (ViT) (Dosovitskiy, 2020), SigLIP (Zhai et al., 2023), or CLIP (Radford et al., 2021)—with the LLM backbone. This alignment is facilitated through linear layers that connect the vision encoder to the backbone, enabling the integration of visual and textual information. For instance, BLIP-2 (Li et al., 2023b) uses OPT (Zhang et al., 2022) and FLAN-T5 (Chung et al., 2022) as backbones, MiniGPT-4 (Zhu et al., 2023) utilizes Vicuna (Chiang et al., 2023), and Qwen-2-VL (Wang et al., 2024) employs Qwen-2-1.5B (Yang et al., 2024) as the language backbone. In this paper, we focus on creating a high-quality multi-image reasoning dataset for instruction tuning instead of large-scale interleaved pretraining datasets like OBELICS (Laurençon et al., 2024a), MINT-1T (Awadalla et al., 2024), and LAION-5B (Schuhmann et al., 2022).
Multi-Image Reasoning Data
Recent advancements in multi-image reasoning instruction tuning datasets include MANTIS (Jiang et al., 2024) and MMDU-45K (Liu et al., 2024f), both aiming to improve reasoning capabilities in VLMs. However, these datasets have limitations in their approaches. MANTIS randomly concatenates single image pairs from LLaVA-665k (Liu et al., 2024a), which often results in uncorrelated images within multi-image scenarios, potentially undermining the complexity of reasoning tasks. MMDU-45K attempts to address this issue by utilizing sentence transformers (Reimers, 2019) with description text and clustering techniques to group related images, but does not consider visual components. The dataset is then further enhanced, assisted by GPT-4 to generate comprehensive answers for the grouped images. Building upon these efforts, SMiR introduces a novel approach that leverages both vital visual and caption information to ensure highly correlated images within multi-image sets with the use of open-source LLMs. These scalable methods leads to the generation of more challenging questions that require deeper analysis and understanding of visual relationships, pushing the boundaries of multi-image reasoning capabilities in VLMs.
Multi-Image Reasoning Benchmarks
Recent VLM benchmarks (Chiang et al., 2024; Lin et al., 2024; Liu et al., 2024d) have made strides by incorporating free-response evaluations, marking a significant improvement over traditional multiple-choice formats. However, these benchmarks still lack a comprehensive approach that combines automatic, multi-turn, and pairwise evaluation capabilities. Our benchmark addresses this gap, drawing inspiration from Auto-Hard-Auto v0.1 (Li et al., 2024b). We have adapted and expanded this framework to enable robust multimodal evaluation, providing a more holistic assessment of VLM performance across complex, multi-image reasoning tasks. This approach allows for a deeper analysis of both the final answers and the underlying reasoning processes employed by VLMs in real-world SMiR-Bench scenarios.
@@ Appendix A Related Works
Table 5: SMiR-Bench scores with GPT-4-Turbo baseline.
@@
SMiR-8B-siglip-llama3 5.0 +.8% (-2.2, 2.7) 183
@@
Appendix B Algorithm Details
B.1 Greedy Cluster Matching Algorithm
@@
B.2 Random Sampling with Iteration
Figure 4: Sampling based on distances between multimodal embeddings of image-caption pairs.
@@ 4. Return $S$
Appendix C Data Samples
@@
C.1 Greedy Cluster Matching Algorithm
Greedy Cluster Matching
1) Orders clusters by size in descending order
This process continues until all clusters are matched or one set is exhausted, demonstrating high efficacy in producing quality image-caption pairs through dual-embedding validation. The score function, which measures overlap normalized by average cluster size, ensures each cluster corresponds to a semantically similar cluster from the other model. The detailed steps of this algorithm and pseudocode are presented in Algorithm 1 (Section B.1).
While this approach effectively leverages both embedding models to confirm spatial relationships and associated semantic meanings, with images matched within clusters being corroborated by two independent embedding models, it has a notable limitation: sampling confined to matched clusters can lead to overly specialized image subjects, such as clustering only sheep-related images rather than diverse animal scenes, as demonstrated in Fig. 5 (Section C.1). This specialization occurs because sampling is restricted to a single matched cluster, potentially limiting the diversity of selected images.
@@ C.1 Greedy Cluster Matching Algorithm
C.2 Random Sampling with Iteration
@@ Appendix D Prompt
D.1 LLaVA Style Prompt